# Boron Fullerenes: From Theoretical Predictions to Experimental Reality

### v1

@@ Abstract
We present a comprehensive first-principles investigation of boron fullerenes and two-dimensional boron sheets, unified under a coordination-based framework. By classifying over a dozen boron nanostructures, including B12, B40, B65, B80, and B92, according to their local atomic environments (4-, 5-, and 6-fold coordination), we identify clear trends in structural stability, electronic properties, and magnetism. A universal energetic scaling relation (with ) captures the convergence of fullerene cohesive energies toward those of 2D boron phases. Notably, we establish one-to-one structural correspondences between select cages and experimentally accessible borophenes: B40 mirrors the -sheet, B65 the -sheet, B80 the -sheet, and B92 the -sheet. These clusters also exhibit large HOMO-LUMO gaps (e.g., eV for B40, 1.14 eV for B92), contrasting with the metallicity of their 2D counterparts and, in the case of B65, spontaneous spin polarization (). Our findings provide a predictive strategy for designing boron nanostructures by leveraging coordination fingerprints, and are further validated by the recent experimental synthesis of the B80 cage. This work bridges zero- and two-dimensional boron chemistry, offering a roadmap for the future synthesis and application of boron-based materials.
@@
The present model is based on the premise that local atomic coordination serves as a chemically meaningful descriptor in electron-deficient systems, such as boron, where conventional or hybridization schemes fail to capture bonding behavior fully [4]. By categorizing each atom within a given cluster as 4-, 5-, or 6-fold coordinated, we establish a coordination-based classification scheme applicable to both finite cages and two-dimensional boron sheets. This approach enables the systematic comparison of structurally diverse fullerenes according to a common, physically interpretable metric. The set of boron hollow cages that was used to establish a new structural classification of these structures is listed in Table 1. Most of the structures listed there have been previously reported in the literature, experimentally (B40 and B80) and theoretically (B12, B60, B65, B80, and B92). The B12 cluster is the building block of all known bulk boron phases. The B60 cage adopts the geometry of the snub dodecahedron. B65 is based on a -symmetric C50 carbon fullerene scaffold reinforced by atoms atop hexagons, and B92 is obtained by extending the well-known B80 cage through the addition of atoms on pentagonal faces. The list is complemented with four additional structures. Based on our previous experience [2] in the construction of B80 from a truncated icosahedron (B60) reinforced with 20 atoms positioned above hexagonal faces, we explored other Archimedean solids as structural templates for hollow boron cages. This search aimed to identify representative structures suitable for our classification framework rather than exhaustively searching for all possible clusters. To obtain the cages, additional boron atoms were placed on select polygonal faces–hexagons or pentagons–to reinforce the cages and emulate bonding patterns observed in the rest of the structures. The B16 cluster is derived from a truncated tetrahedron with four capping atoms on hexagons, while B32 and B42 originate from a truncated octahedron (with eight capping atoms on hexagons) and an icosidodecahedron (with twelve capping atoms on pentagons), respectively. Finally, the B24 cluster adopts the geometry of the snub cube. The full list of structures considered, including their symmetry, coordination environments, and references to prior studies, is summarized in Table 1. To provide a unified energetic comparison across clusters of different sizes, we employ a size-dependent scaling law for the cohesive energy per atom of the form
@@
| 5-coordinated atoms | |
| B24 | B60 |
| 5,6-coordinated atoms | |
| B16 | B32 |
| 5,6-coordinated atoms (Boustani cages) | |
| B12 | B42 |
| 4,5-coordinated atoms | |
| B40 | -sheet |
| 4,5,6-coordinated atoms | |
| B65 | -sheet |
| 5,6-coordinated atoms | |
| B80 | -sheet |
| 5,6-coordinated atoms (Boustani structures) | |
| B92 | -sheet |
@@
This coordination-based framework enables a consistent structural classification of boron fullerenes and provides a direct link to the local bonding motifs found in experimentally and theoretically known borophene sheets. Its predictive value lies in establishing an organizational principle that bridges the structural diversity of boron clusters with the topology of extended 2D phases. The local atomic coordination environment strongly modulates the relative energetic stability of boron fullerenes. Figure 1 displays the energy difference per atom, , relative to the cohesive energy of B80 (), plotted as a function of the total number of atoms . Structures are grouped according to dominant coordination motifs–purely 5-fold, mixed 5,6-fold, hybrid 4,5,6-fold, and so forth–and fitted to the empirical scaling law . The fitted exponent was determined for families of clusters composed of purely 5-fold and mixed 5,6-fold coordinated boron atoms, yielding values close to . For clusters with insufficient data–namely, those with 4,5-fold and 4,5,6-fold coordination– was fixed at 0.9 and only the prefactor was fitted. This approach captures the general trend that strain energy per atom decreases with increasing cluster size, as larger fullerenes approximate a spherical geometry and reduce distortions from planar bonding preferences. The most stable fullerenes within each coordination family consistently exhibit a higher degree of atomic coordination. For instance, B80 and B92 (with 20 and 80 atoms in 6-fold coordination, respectively) have cohesive energies of 5.813 eV and 5.781 eV per atom, among the highest in the set. B92 is derived from B80 by capping its pentagonal faces with twelve additional atoms; while it features more 6-fold coordinated atoms, B80 remains slightly more stable due to a better balance between 5- and 6-fold coordinated sites. In contrast, purely 5-coordinated structures such as B24 and B60 are less favorable, with eV and 5.671 eV, respectively, despite their high symmetry ( and ). Structures with mixed coordination, such as B40 (, ) and B65 (, , ), display enhanced energetic stability ( eV and 5.747 eV, respectively) highlighting the advantage of combining multiple coordination environments. These trends support the conclusion that coordination number distributions are predictive of fullerene stability, providing a quantitative metric for comparing structurally diverse boron nanoclusters.
@@
Visual analysis of optimized cage geometries (Figures 2 and 3) reveals that local coordination motifs not only determine energetic preferences but also provide a structural bridge between 0D fullerenes and 2D boron sheets. B24 and B60, although purely 5-coordinated and energetically less favorable despite their high symmetry, are structurally related to the periodic snub-sheet, highlighting their role as finite analogues of extended 2D phases. More complex coordination patterns yield deeper structural correspondences: B40 shares its 4,5-coordination scheme with the -sheet and adopts symmetry; B65 reflects the atomic environment of the -sheet with a mixed 4,5,6-fold coordination; B80, with atoms centered on hexagons, mirrors the -sheet; and B92 further extends this motif by adding twelve atoms atop the pentagons of B80, closely resembling the densely packed -sheet. Overall, structures combining diverse coordination environments are preferred over uniformly coordinated ones–both in fullerenes and borophenes–and the observed structural analogies provide a rational framework for interpreting boron fullerenes as finite reconstructions of stable 2D boron motifs.
@@
Table 1 details the electronic gaps and total magnetic moments for all clusters and sheet models. Clusters with substantial 6-fold coordination, such as B80 and B92, possess large HOMO–LUMO gaps (1.02 eV and 1.14 eV, respectively), indicative of electronic stability and low chemical reactivity. Notably, B40 exhibits the largest gap (1.78 eV) despite its lower coordination, likely due to enhanced orbital localization arising from its compact and symmetric geometry. In contrast, purely 5-fold coordinated clusters such as B24 and B60 have much smaller gaps (0.31 eV and 0.13 eV), approaching metallic behavior. These trends underscore the role of the coordination environment in shaping electronic properties, with higher and mixed coordination leading to more stable electronic configurations. Unlike the cages, their corresponding two-dimensional boron sheets are metallic under the same computational conditions, highlighting the distinct electronic character that emerges in finite versus extended systems. Magnetism arises in clusters with mixed or frustrated coordination patterns: B65 exhibits a net magnetic moment of , absent in its 2D analogue; similar spin polarization is found in B12 (), B16 (), and B42 (). None of the extended boron sheets–, , , or –exhibit magnetization, suggesting that finite size and geometric frustration are essential for inducing magnetic moments. These results demonstrate that both electronic and magnetic properties depend sensitively on local coordination and its spatial distribution, providing a predictive handle for designing boron nanostructures with targeted functionalities.
@@
The observed trends in cohesive energy, electronic structure, and magnetism reinforce the central premise of this work: that local atomic coordination serves as a chemically meaningful and predictive descriptor for boron nanostructures. Clusters with coordination fingerprints similar to known borophene phases tend to exhibit comparable energetic and electronic behavior, while purely 5-coordinated or irregular topologies deviate from this trend. As shown in Figures 1–3 and Table 1, this coordination-driven classification enables the evaluation of structurally and electronically diverse fullerenes on a common footing, offering a systematic framework for interpreting their stability, electronic properties, and magnetism. This approach provides a practical set of design rules for constructing new boron-based nanostructures with tailored functionalities, leveraging the predictive power of local bonding environments.
@@
Summarizing, we have presented a comprehensive coordination-based classification of boron fullerenes, linking their energetic, electronic, and magnetic properties to their local bonding environments. Using first-principles calculations, we analyzed a representative set of cage geometries, ranging from purely 5-coordinated high-symmetry clusters to hybrid 4,5,6-coordinated cages with direct analogues in borophene sheets. Energetic stability is found to increase with coordination number, with the most stable structures–B80 and B92–featuring large fractions of 6-fold coordinated atoms. The relative energy scaling follows a universal trend with cluster size, well described by with , indicating convergence toward sheet-like cohesive energies. Clusters with significant 6-fold coordination, such as B80 and B92, exhibit large HOMO–LUMO gaps (1.02 and 1.14 eV), indicating enhanced electronic stability. B40 shows the largest gap (1.78 eV) despite lower coordination, due to orbital localization from its compact geometry. In contrast, purely 5-fold coordinated clusters like B24 and B60 have much smaller gaps (0.31 and 0.13 eV), approaching metallic behavior. The corresponding 2D boron sheets are metallic, underscoring the electronic differences between finite and extended systems. Our coordination-based framework also identifies one-to-one correspondences between fullerene cages and known borophene sheets. B40 mirrors the -sheet, B80 mirrors the -sheet, B65 corresponds to the -sheet, and B92 closely resembles the -sheet. According to this simple model, the experimentally accessible structures (B40, -sheet) and (B80, -sheet) can be complemented with two new pairs–(B65, -sheet) and (B92, -sheet)–for which the cage clusters should also, in principle, be experimentally feasible. These findings demonstrate that local coordination provides a unifying and predictive metric for organizing and designing boron nanostructures across dimensions. The structural and electronic coherence between fullerenes and borophenes revealed here offers a roadmap for identifying stable, functional boron materials suitable for experimental realization.

### v2

# Coordination-Driven Classification and Energetic Scaling of Boron Fullerenes and Borophenes

@@ Abstract
We present a comprehensive first-principles investigation of boron fullerenes and two-dimensional boron sheets, unified under a coordination-based framework. By classifying over a dozen boron nanostructures, including B12, B40, B38, B65, B80, and B92, according to their local atomic environments (4-, 5-, and 6-fold coordination), we identify clear trends in structural stability, electronic properties, and magnetism. A universal energetic scaling relation , with or depending on the coordination family, captures the convergence of fullerene cohesive energies toward those of 2D boron phases. Notably, we establish one-to-one structural correspondences between select cages and experimentally accessible borophenes: B40 mirrors the -sheet, B65 the -sheet, B80 the -sheet, and B92 the -sheet. Our analysis reveals two distinct families of boron nanostructures based on the scaling exponent: one with , comprising structures with significant 6-fold coordination, and another with , which includes B40, B38, and two experimentally observed borophenes. These clusters also exhibit large HOMO–LUMO gaps (e.g., eV for B40, 1.14 eV for B92), contrasting with the metallicity of their 2D counterparts and, in the case of B65, spontaneous spin polarization (). Our findings provide a predictive strategy for designing boron nanostructures by leveraging coordination fingerprints, and are further validated by the recent experimental synthesis of the B80 cage. This work bridges zero- and two-dimensional boron chemistry, offering a roadmap for the future synthesis and application of boron-based materials.
@@
The present model is based on the premise that local atomic coordination serves as a chemically meaningful descriptor in electron-deficient systems, such as boron, where conventional or hybridization schemes fail to capture bonding behavior fully [4]. By categorizing each atom within a given cluster as 4-, 5-, or 6-fold coordinated, we establish a coordination-based classification scheme applicable to both finite cages and two-dimensional boron sheets. This approach enables the systematic comparison of structurally diverse fullerenes according to a common, physically interpretable metric. The set of boron hollow cages that was used to establish a new structural classification of these structures is listed in Table 1. Most of the structures listed there have been previously reported in the literature, experimentally (B40 and B80) and theoretically (B12, B38, B60, B65, B80, and B92). The B12 cluster is the building block of all known bulk boron phases. The B38 cluster is a fullerene-like structure featuring four hexagons and 56 triangles in a highly symmetric cage, and it exhibits notable energetic and chemical stability, including a large HOMO–LUMO gap and double aromaticity. The B60 cage adopts the geometry of the snub dodecahedron. B65 is based on a -symmetric C50 carbon fullerene scaffold reinforced by atoms atop hexagons, and B92 is obtained by extending the well-known B80 cage through the addition of atoms on pentagonal faces. The list is complemented with four additional structures. Based on our previous experience [2] in the construction of B80 from a truncated icosahedron (B60) reinforced with 20 atoms positioned above hexagonal faces, we explored other Archimedean solids as structural templates for hollow boron cages. This search aimed to identify representative structures suitable for our classification framework rather than exhaustively searching for all possible clusters. To obtain the cages, additional boron atoms were placed on select polygonal faces–hexagons or pentagons–to reinforce the cages and emulate bonding patterns observed in the rest of the structures. The B16 cluster is derived from a truncated tetrahedron with four capping atoms on hexagons, while B32 and B42 originate from a truncated octahedron (with eight capping atoms on hexagons) and an icosidodecahedron (with twelve capping atoms on pentagons), respectively. Finally, the B24 cluster adopts the geometry of the snub cube. The full list of structures considered, including their symmetry, coordination environments, and references to prior studies, is summarized in Table 1. To provide a unified energetic comparison across clusters of different sizes, we employ a size-dependent scaling law for the cohesive energy per atom of the form
@@
| 5-coordinated atoms | ||||||||
| 5,6-coordinated atoms | ||||||||
| 5,6-coordinated atoms (Boustani cages) | ||||||||
| 4,5,6-coordinated structures | ||||||||
| 4,5-coordinated structures | ||||||||
@@
This coordination-based framework enables a consistent structural classification of boron fullerenes and provides a direct link to the local bonding motifs found in experimentally and theoretically known borophene sheets. Its predictive value lies in establishing an organizational principle that bridges the structural diversity of boron clusters with the topology of extended 2D phases. The local atomic coordination environment strongly modulates the relative energetic stability of boron fullerenes. Figure 1 displays the energy difference per atom, , relative to the cohesive energy of B80 (), plotted as a function of the total number of atoms . Structures are grouped according to dominant coordination motifs—purely 5-fold, mixed 5,6-fold, hybrid 4,5,6-fold, and so forth—and fitted to the empirical scaling law . The fitted exponent distinguishes two coordination families: one with for clusters dominated by 5- and 5,6-fold coordination, and another with for structures involving 4,5- and 4,5,6-fold motifs. This approach captures the general trend that strain energy per atom decreases with increasing cluster size, as larger fullerenes approximate a spherical geometry and reduce distortions from planar bonding preferences. The most stable fullerenes within each coordination family consistently exhibit a higher degree of atomic coordination. For instance, B80 and B92 (with 20 and 80 atoms in 6-fold coordination, respectively) have cohesive energies of 5.813 eV and 5.781 eV per atom, among the highest in the set. B92 is derived from B80 by capping its pentagonal faces with twelve additional atoms. At the same time, it features more 6-fold coordinated atoms, B80 remains slightly more stable due to a better balance between 5- and 6-fold coordinated sites. In contrast, purely 5-coordinated structures such as B24 and B60 are less favorable, with eV and 5.671 eV, respectively, despite their high symmetry ( and ). Structures with mixed coordination, such as B40 (, ), B38 (, , ), and B65 (, , ), display enhanced energetic stability ( eV, 5.703 eV, and 5.747 eV, respectively), highlighting the advantage of combining multiple coordination environments. These trends support the conclusion that coordination number distributions are predictive of fullerene stability, providing a quantitative metric for comparing structurally diverse boron nanoclusters.
@@
Visual analysis of optimized cage geometries (Figure 2) reveals that local coordination motifs not only determine energetic preferences but also provide a structural bridge between 0D fullerenes and 2D boron sheets. B24 and B60, although purely 5-coordinated and energetically less favorable despite their high symmetry, are structurally related to the periodic snub-sheet, highlighting their role as finite analogues of extended 2D phases. More complex coordination patterns yield deeper structural correspondences: B40 shares its 4,5-coordination scheme with the -sheet and adopts symmetry; B65 reflects the atomic environment of the -sheet with a mixed 4,5,6-fold coordination; B80, with atoms centered on hexagons, mirrors the -sheet; and B92 further extends this motif by adding twelve atoms atop the pentagons of B80, closely resembling the densely packed -sheet. Overall, structures combining diverse coordination environments are preferred over uniformly coordinated ones–both in fullerenes and borophenes–and the observed structural analogies provide a rational framework for interpreting boron fullerenes as finite reconstructions of stable 2D boron motifs.
@@
Table 1 details the electronic gaps and total magnetic moments for all clusters and sheet models. Clusters with substantial 6-fold coordination, such as B80 and B92, possess large HOMO–LUMO gaps (1.02 eV and 1.14 eV, respectively), indicative of electronic stability and low chemical reactivity. Notably, B40 exhibits the largest gap (1.78 eV) despite its lower coordination, likely due to enhanced orbital localization arising from its compact and symmetric geometry. Other structures with mixed coordination, including B38 and B65, also show moderate-to-large gaps (1.12 eV and 0.17 eV, respectively), underscoring the stabilizing role of hybrid bonding motifs. In contrast, purely 5-fold coordinated clusters such as B24 and B60 have much smaller gaps (0.31 eV and 0.13 eV), approaching metallic behavior. These trends underscore the role of the coordination environment in shaping electronic properties, with higher and mixed coordination leading to more stable electronic configurations. Unlike the cages, their corresponding two-dimensional boron sheets are metallic under the same computational conditions, highlighting the distinct electronic character that emerges in finite versus extended systems. Magnetism arises in clusters with mixed or frustrated coordination patterns: B65 exhibits a net magnetic moment of , absent in its 2D analogue; similar spin polarization is found in B12 (), B16 (), and B42 (). None of the extended boron sheets—, , , or —exhibit magnetization, suggesting that finite size and geometric frustration are essential for inducing magnetic moments. These results demonstrate that both electronic and magnetic properties depend sensitively on local coordination and its spatial distribution, providing a predictive handle for designing boron nanostructures with targeted functionalities.
@@
The observed trends in cohesive energy, electronic structure, and magnetism reinforce the central premise of this work: that local atomic coordination serves as a chemically meaningful and predictive descriptor for boron nanostructures. Clusters with coordination fingerprints similar to known borophene phases tend to exhibit comparable energetic and electronic behavior, while purely 5-coordinated or irregular topologies deviate from this trend. As shown in Figures 1 and 2, and Table 1, this coordination-driven classification enables the evaluation of structurally and electronically diverse fullerenes on a common footing, offering a systematic framework for interpreting their stability, electronic properties, and magnetism. This approach offers a practical set of design rules for constructing new boron-based nanostructures with tailored functionalities, leveraging the predictive power of local bonding environments.
@@
Summarizing, we have presented a comprehensive coordination-based classification of boron fullerenes, linking their energetic, electronic, and magnetic properties to their local bonding environments. Using first-principles calculations, we analyzed a representative set of cage geometries, ranging from purely 5-coordinated high-symmetry clusters to hybrid 4,5,6-coordinated cages with direct analogs in borophene sheets. Energetic stability is found to increase with coordination number, with the most stable structures–B80 and B92–featuring large fractions of 6-fold coordinated atoms. The relative energy scaling follows two distinct universal trends with cluster size, well described by with or , indicating convergence toward sheet-like cohesive energies depending on coordination type. Clusters with significant 6-fold coordination, such as B80 and B92, exhibit large HOMO–LUMO gaps (1.02 and 1.14 eV), indicating enhanced electronic stability. B40 shows the largest gap (1.78 eV) despite lower coordination due to orbital localization from its compact geometry. In contrast, purely 5-fold coordinated clusters like B24 and B60 have much smaller gaps (0.31 and 0.13 eV), approaching metallic behavior. The corresponding 2D boron sheets are metallic, underscoring the electronic differences between finite and extended systems. Our coordination-based framework also identifies one-to-one correspondences between fullerene cages and known borophene sheets. B40 mirrors the -sheet, B80 mirrors the -sheet, B65 corresponds to the -sheet, and B92 closely resembles the -sheet. According to this simple model, the experimentally accessible structures (B40, -sheet) and (B80, -sheet) can be complemented with two new pairs–(B65, -sheet) and (B92, -sheet)–for which the cage clusters should also, in principle, be experimentally feasible. These findings demonstrate that local coordination provides a unifying and predictive metric for organizing and designing boron nanostructures across dimensions. The structural and electronic coherence between fullerenes and borophenes revealed here offers a roadmap for identifying stable, functional boron materials suitable for experimental realization.